To buy both a new car and a new house, Tina sought two loans totaling $23,067. The simple interest rate on the first loan was 6.8%, while the simple interest rate on the second loan was 7.7%. At the end of the first year, Tina paid a combined interest payment of $1765.70. What were the amounts of the two loans?
Answer by stanbon(75887)

To buy both a new car and a new house, Tina sought two loans totaling $23,067. The simple interest rate on the first loan was 6.8%, while the simple interest rate on the second loan was 7.7%. At the end of the first year, Tina paid a combined interest payment of $1765.70. What were the amounts of the two loans?
---------
Equation:
int + int = 1765.70
----
0.068x + 0.077(23067-x) = 1765.70
-----
Multiply thru by 1000
68x + 77*23067 - 77x = 1765700
-9x = -10459
x = $1162.11(amt. invested at 6.8%)
---------------------
23,067-x = $21904.89 (amt. invested 7.7%)
===============================
Cheers,
Stan H.
================